Two marbles are drawn (without replacement) from a bag containing 3 green, 6 yellow,and 4 red marbles.
What is the probability that the marbles are
of different colors?

Answer by Positive_EV(69)

The event that the marbles are different colors is the complement of the event that the marbles are the same color. Thus, calculate the probability that the marbles are the same color, then subtract this probability from 1 to find the probability they are different colors.

P(2 green) = (3/13)(2/12) = 1/26
P(2 yellow) = (6/13)(5/12) = 5/26
P(2 red) = (4/13)(3/12) = 2/26 = 1/13

P(2 of the same) = P(2 green) + P(2 yellow) + P(2 red) = 1/26 + 5/26 + 2/26 = 8/26 = 4/13.
P(2 different) = 1 - P(2 of the same) = 1 - 4/13 = 9/13.